Tuesday, May 3, 2016
03 May 2016 | 17:32 PM - We're acknowledging an issue with the routing of our 22.214.171.124/24 IP block and have engaged our upstream carriers to rectify the issue. If your IP begins with 104.251.212, then you are effected by this. We'll update this announcement as we have more information. 03 May 2016 | 17:55 PM - This has been resolved.
Sunday, May 1, 2016
We are aware that ukpure11 is down at the moment. We are troubleshooting it.
Thursday, April 28, 2016
We are currently doing a control panel maintenance. If you see "Unknown Error" on the VPS details page, this is a known issue. We will update this announcement once this is done.
Monday, April 11, 2016
11 April 2016 | 08:51 AM - the following advance notice was sent of outage: Hello, We are writing to inform you of an upcoming short outage on our ukpure10 server where you have active service. We will be shutting this server down at approximately 21:45 Monday, April 11, 2016 to swap RAM within the server. We have this action planned for a 22:00 start time with the...
Saturday, April 2, 2016
As scheduled and notified, the UK hardware maintenance is in progress now. Your would see around 30 mins of downtime. Please do not open a ticket and follow this announcement for updates. Update 09:12 PM CST: We are still waiting for an update for node ukpure6 to come back up. We are also going to investigate the cause of the lengthy downtime of this first node and prevent this to happen on...
Friday, April 1, 2016
Due to soft lock ups, we are rebooting dalpure6. Please wait till your VPS comes back online. Update 05:46 AM: The VPS's are booting back up. Everything should come back up in the following 10 minutes. Update 05:55 AM: Everything is back up.
Thursday, March 31, 2016
31 March 2016 | 08:17 AM - we are investigating a connectivity issue with our ukpure11 server. 31 March 2016 | 08:54 AM - we rectified this situation by rebooting the server to clear the load lock. The server is back online and VMs are booting now. Your VM should be online soon if it is not already. 31 March 2016 | 09:16 AM - the datacenter rebooted the host node...
Monday, March 21, 2016
21 March 2016 | 17:15 PM - Acknowledging high packet loss for Dallas storage services as of a few minutes ago, currently investigating. 21 March 2016 | 17:31 PM - This has been rectified and all services are back online.
Saturday, March 19, 2016
19 March 2016 | 21:34 PM - Further to the email you received, the daldb migration is in progress now, slightly past schedule. Our apologies for the delay. For your reference here is the email we sent you: http://disclosed.info/?b46ba2c322fb4429#yGflzHchyNQtJW//qnQoKNgCwFJ/2Ii1UTZeEn4fS6E= 19 March 2016 | 22:18 PM - The transfer continues to be in progress. 19 March 2016 |...
Saturday, March 12, 2016
There is an issue requiring the reboot of this now. Vms will be going down for a few minutes Update: VMs should now be coming back online